I’m hooked on YouTube physics videos.

If E= MC2, how do you explain that since C2 is impossible,how does this make sense?


1
Expert's answer
2021-04-14T10:47:16-0400

You probably mean that no body, that has mass, can move with the velocity v=c. The answer to your question lies in understanding what "E=mc^2" indeed means. What is this energy? It is not just some energy. In general, there are different kinds of energy.

Let's imagine a person in the gym located on the 3rd floor. This person is running on the running track with speed v. So the kinetic energy is "mv^2\/2". This is the energy his/her muscles develop to run with the speed v. This person also has potential energy, because he/she is on the 3rd floor. It the person jumps from the window on the ground, the potential energy will change, and the difference "mg(h_2-h_1)" will transfer into the energy of deformation (this energy can, for example, break bones or do another harmful things to the body). And of course this person cannot move with v=c. Then where do we use "E=mc^2"? Seems like we have already described the energy of this person. "E = mv^2\/2 + mgh_2".

Not exactly. "E=mc^2" is the store of energy that every massive body has. If you put together electron and positron, particle and antiparticle, they will annihilate - their mass will be converted to pure energy. From this "pure energy" other particles can be born, but this is another, more complicated story. So if we take person and exactly the same anti-person, they will annihilate too. Anti-person will be made of anti-matter (all particles are substituted by their anti particles). How much energy will be released? And this is the answer. "E=mc^2". We do not use this energy in mechanics, because we do not destroy mechanical bodies. To get "E=mc^2" we need to absolutely destroy the body. This energy is important for particles though, because particles can interact with other particles, they can decay or born new particles. In all these process "E=mc^2" is important. Particles are more likely to destroy themselves fully (or more scientifically, to convert their rest mass to energy).